suppose that you borrow $60,000 at 9% compounded monthly over five years. Knowing that the 9% represents the market interest rate. you compute the monthly payment in actual dollars as $1,245.51. If the average monthly general inflation rate is expected to be 0.25%, determine the equivalent equal monthly payment series in constant dollars. 7,
Answer #1

Equivalent equal monthly payment series in constant dollars

= Actual dollars (1+0.0025)-60 = 1245.51(1+0.0025)-60 = 1245.51 * 0.861 = $1072.38